OpenStudy (anonymous):

how many roots does the equation z^4 + 16z* = 0 have? z * is the complex conjugate

OpenStudy (anonymous):

Multiply z on both sides z^4 + 16|z|^2 /z =0 Six Solutions

OpenStudy (anonymous):

OK one of the roots is z=0. Now for the nonzero roots I do understand your new equation form. However, how do you conclude 5 more roots from z^4 + 16|z|^2/z = 0?

OpenStudy (anonymous):

\[\frac{z^5+16|z|^2}{z}=0\] get it

OpenStudy (anonymous):

YOu are in 11th grade preparing for JEE

OpenStudy (anonymous):

They include complex roots